# Theory of Mind in Large Language Models: Assessment and Enhancement

### v1

@@ Abstract
Theory of Mind (ToM)—the ability to infer and reason about others’ mental states—is fundamental to human social intelligence. As Large Language Models (LLMs) become increasingly integrated into daily life, it is crucial to assess and enhance their capacity to interpret and respond to human mental states. In this paper, we review LLMs’ ToM capabilities by examining both evaluation benchmarks and the strategies designed to improve them. We focus on widely adopted story-based benchmarks and provide an in-depth analysis of methods aimed at enhancing ToM in LLMs. Furthermore, we outline promising future research directions informed by recent benchmarks and state-of-the-art approaches. Our survey serves as a valuable resource for researchers interested in advancing LLMs’ ToM capabilities.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have achieved remarkable success across various tasks, particularly in Natural Language Processing (NLP) Qin et al. (2023). Currently, there is growing interest to explore the cognitive abilities of LLMs through novel and intriguing tasks, including assessments of their Emotional Intelligence (EI) Yang et al. (2023); Huang et al. (2024); Amin et al. (2024); Sabour et al. (2024); Liu et al. (2024), strategic reasoning capabilities Gandhi et al. (2023b); Zhang et al. (2024b), and ToM capabilities Wu et al. (2023); van Duijn et al. (2023); Wu et al. (2024); Shapira et al. (2024); Chen et al. (2024); Jin et al. (2024). Among these capabilities, there is an ongoing debate about whether LLMs truly possess ToM abilities. Some studies suggest that LLMs demonstrate promising signs of ToM competence Bubeck et al. (2023); Kosinski (2024); Street et al. (2024), while others contend that these abilities are often superficial and unstable Shapira et al. (2024); Ullman (2023); Ma et al. (2023a); Verma et al. (2024); Zhu et al. (2024).
Figure 1: Scope of Our Survey. This paper examines both the evaluation and enhancement of theory of mind capabilities in LLMs. For evaluation, we cover passive, story-based benchmarks and adaptable interactive benchmarks. For enhancement, we discuss recent effective methods, including prompt-only approaches and strategies incorporating additional techniques, as well as earlier methods from the era of small language models. Due to space constraints, interactive benchmarks and earlier methods (indicated in the dashed area) are provided in the Appendix E and Appendix F.
@@
Consequently, our work focuses on newly proposed story-based benchmarks from 2023–2024 that have not been covered in existing reviews, examining their evolution. We also provide a detailed review of the recent approaches aimed at enhancing LLMs’ ToM capabilities. As far as we know, this is the first study to offer such an in-depth analysis of both the evaluation and enhancement of LLMs’ ToM capabilities.
As illustrated in Figure 1, our survey encompasses both the evaluation of LLMs’ ToM capabilities—highlighting widely used story-based benchmarks—and the strategies employed to enhance their performance, with the aim of establishing a foundation for future research in this area. Our contributions are summarized as follows:
  • Broad Survey. This paper provides the first broad survey that addresses both the evaluation and enhancement of LLMs’ ToM capabilities. It offers an in-depth analysis of recently proposed story-based benchmarks Ma et al. (2023b) and the strategies developed to improve LLM performance in this domain.
@@ 2 Theory of Mind
The term "orders" refers to the number of mental state attributions required to answer a given question Wu et al. (2023). For example, the question "Where will Sally look for her marble?" is considered first-order, while "Where does Anne think Sally will look for her marble?" is classified as second-order. When Sally’s belief is incorrect, the question is categorized as a false-belief question; conversely, if her belief is accurate, it is referred to as a true-belief question.
@@
As shown in Table 1, more than ten benchmarks have been recently proposed to assess the ToM capabilities of LLMs. The majority of these benchmarks are based on well-established psychological tests, including the Sally-Anne Test Baron-Cohen et al. (1985) and the Smarties Task Gopnik and Astington (1988), both of which are discussed in detail in Appendix A.
@@ 3 Benchmarks for Evaluating ToM Capabilities in LLMs
3.1 Text-Only Benchmark
@@ 3.1.1 Overview of Text-Only Benchmarks
TOMCHALLENGES Ma et al. (2023a). Ma et al. (2023a) identified that inconsistencies in evaluating LLMs’ ToM capabilities might arise from variations in evaluation methods and prompts. To address this, they designed six distinct questions for each "order" of ToM reasoning about each narrative story, using templates that span three generation settings: fully-constrained, semi-constrained, and open-ended. The stories are adapted from classic tasks like the Sally-Anne and Smarties tasks Baron-Cohen et al. (1985); Gopnik and Astington (1988), covering up to second-order belief questions. To enhance grading efficiency, they proposed an auto-grader based on the GPT-4222 https://platform.openai.com/docs/models/gp model. However, the effectiveness of this automatic evaluation tool may depend on the model’s own ToM capabilities.
@@
NegotiationToM Chan et al. (2024). Grounded in the Belief-Desire-Intention (BDI) agent modeling theory Bratman (1987), NegotiationToM is designed to rigorously evaluate machine ToM in real-world negotiation scenarios featuring multiple rounds of dialogue. It assesses a range of multidimensional mental states, including desires, second-order beliefs, and intentions. NegotiationToM is based on the CaSiNo dataset Chawla et al. (2021), which comprises negotiation scenarios between two parties. Although it is the first ToM benchmark in negotiation contexts, it remains a passive benchmark. Additionally, its reliance on an existing dataset introduces a potential risk of data contamination.
@@ 3.1.2 Trends in Text-Only Benchmarks
  • Questions. To enable robust evaluation, benchmark questions have become more diverse and complex, evolving from multiple-choice question answering Le et al. (2019) to open-ended question answering Ma et al. (2023a); Kim et al. (2023). Additionally, the content of the questions has changed, for instance, asking about the beliefs of a group of people rather than a single individual Sileo and Lernould (2023); Kim et al. (2023).
  • Mental States. Benchmarks have expanded from evaluating only beliefs Le et al. (2019) to covering a wider range of mental states Sileo and Lernould (2023); Chan et al. (2024); Chen et al. (2024), including percepts, desires, and even social commonsense knowledge Xu et al. (2024a).
3.2 Multimodal Benchmark
In addition to the aforementioned developments, some researchers have recognized that ToM reasoning goes beyond simple text comprehension Chen et al. (2024); Jin et al. (2024); Shi et al. (2024), which has led to the introduction of several multimodal ToM benchmarks.
@@ 3.2 Multimodal Benchmark
MuMA-ToM Shi et al. (2024). Building on videos generated in VirtualHome Puig et al. (2018), this benchmark introduces the first multimodal, multi-agent ToM evaluation focusing on mental reasoning in embodied interactions. It includes three types of questions: belief inference, social goal inference, and belief-of-goal inference (a form of second-order belief). Each question presents three options, with one being the most likely to be correct. Unlike MMToM-QA Jin et al. (2024), MuMA-ToM provides distinctly separate information available through specific modalities, enabling a deeper understanding of how models integrate multimodal inputs to accurately answer each question. The text input is presented in two forms: conversations between two agents in the video or textual descriptions of half of the video. MuMA-ToM is currently limited to three social goals and two agents. Additionally, as it relies on video content from VirtualHome, further efforts are required to enhance its sim-to-real transferability for real-world applications.
@@ 5 Future Directions
Most evaluations indicate that LLMs still lack robust ToM reasoning abilities. A clear direction for future work is to develop strategies to enhance LLMs’ ToM reasoning capabilities. To improve strategies for enhancing LLMs’ ToM capabilities, more benchmarks are required—both for evaluating ToM reasoning and as resources for fine-tuning LLMs. In this section, we first explore common future directions for both benchmarks and strategies, followed by a discussion of directions that are specific to either enhancement or evaluation [^4] Due to space limitations, additional future directions are discussed in the Appendix G..
Common future directions for both benchmarks and strategies involve:
  • Expanding the scope of mental states: As shown in Table 1 and Table 3, current benchmarks and strategies predominantly focus on belief-related reasoning, while other mental states require more attention Ma et al. (2023b). Future benchmarks should explore a broader range of mental states. To better evaluate LLMs’ ToM capabilities, a wider variety of psychological tests should be integrated into the benchmarking process. Fu et al. (2023) compiled a list of 127 tests designed to measure the ToM capabilities of children from birth to 12 years old, which could be further adapted for testing LLMs as the technology advances.
  • Addressing multi-modal ToM reasoning: Humans interact with the world through multiple channels—an aspect that simple text-based stories cannot fully capture Liu et al. (2023). Research on ToM capabilities in LLMs aims to develop models that function as agents with robust ToM skills, potentially achieving human-level competence Ma et al. (2023b). Achieving this goal requires the ability to infer mental states from visual, auditory, contextual, and other cues. For instance, incorporating multimodal content such as short films Dziobek et al. (2006) or cartoons Völlm et al. (2006); Parmar et al. (2024) into evaluation benchmarks could enhance ToM assessments. Moreover, current methods primarily focus on multiple-choice question answering in household contexts, and converting video information to text may lead to the loss of crucial details. Developing strategies that effectively handle multimodal inputs and preserve comprehensive information remains an important direction for future research.
  • Active benchmarks and strategies for agentic decision-making: Passive benchmarks Ma et al. (2023b); Chan et al. (2024) are insufficient. More research is needed to leverage LLMs as agents capable of making decisions in complex environments Ma et al. (2023b); Li et al. (2023); Zhou et al. (2023a, b), enabling deeper investigation of their ToM capabilities.
@@
  • Exploring joint approaches: Most existing methods rely on pipeline architectures with no feedback loops between stages. Error propagation is a significant issue in pipeline approaches Yang and Mitchell (2016); Liu et al. (2018). For example, in SYMBOLICTOM, if the belief graph for a character is incorrect, it becomes difficult to obtain the correct answer when the question relates to this flawed belief graph. However, if conflicts are detected during the inference stage and feedback is provided during the belief graph construction stage, the accuracy of the inference and the coherence of the results can be improved. Investigating joint or iterative approaches that integrate feedback to improve ToM reasoning remains an open research question.
  • Progressive learning strategies for tackling ToM task complexity: Given that ToM tasks can vary significantly in complexity, a continual/curriculum learning Chen and Liu (2018); Soviany et al. (2022); Wang et al. (2024) strategy may be needed to address these challenges progressively, starting with simpler tasks and advancing to more complex ones.
Finally, we believe that evaluating reasoning processes is crucial. Simply assessing the correctness of answers is insufficient, benchmarks should facilitate the evaluation of reasoning processes Kawabata and Sugawara (2023); Jung et al. (2024); Xu et al. (2024b), which calls for the development of more effective and automated evaluation strategies.
@@ 6 Conclusion
Appendix A Psychology Tests
@@ A.1 Sally-Anne Test
  • Reality Question "Where is the marble really?"
  • Memory Question "Where was the marble in the beginning?"
@@ A.3 Smarties Task
The Smarties task Perner et al. (1987), sometimes referred to as "unexpected contents [^5] https://en.wikipedia.org/wiki/Theory_of_mind" or "appearance-reality [^6] http://www.sfu.ca/~hedberg/ToM.pdf" task, involves asking children what they think is inside a box that looks like it contains smarties. After the child guesses "smarties", they are shown that the box actually holds a different item, such as pencils. The box is then resealed, and the child is asked what they think someone else, who hasn’t seen the real contents, will believe is inside. The child succeeds if they say the other person will think "smarties" are in the box, and fails if they say the other person will think it contains pencils. According to Gopnik and Astington’s research, children generally pass this task by the age of four or five Gopnik and Astington (1988).
@@ Appendix C Overview of Additional Text-Only Benchmarks
Benchmarks Mental States
ToMi Le et al. (2019)
MindGames Sileo and Lernould (2023)
HI-TOM Wu et al. (2023)
TOMCHALLENGES Ma et al. (2023a)
FANTOM Kim et al. (2023)
BigToM Gandhi et al. (2023a)
OpenToM Xu et al. (2024a)
NegotiationToM Chan et al. (2024)
TOMBENCH Chen et al. (2024)
COMMON-TOM Soubki et al. (2024)
SimpleToM Gu et al. (2024)
MMToM-QA Jin et al. (2024)
MuMA-ToM Shi et al. (2024)
Table 1: Benchmarks and Their Coverage of Mental States: The mental state categories are derived from the ATOMS Beaudoin et al. (2020). In this paper, we treat the "goals" referenced in MMToM-QA and MuMA-ToM as equivalent to "intentions" in ATOMS. Benchmarks are divided into two categories: text-only benchmark and multimodal benchmark
@@
Figure 3: A comparison of methods for enhancing LLMs’ ToM capabilities through different prompting techniques, using the following narrative as context. Story: Alice and Bob are in a room with a basket and a box. Alice puts some celery in the basket and leaves the room. Bob then moves the celery into the box. Sclar et al. (2023)

### v2

@@ Abstract
Theory of Mind (ToM)—the ability to reason about the mental states of oneself and others—is a cornerstone of human social intelligence. As Large Language Models (LLMs) become increasingly integrated into daily life, understanding their ability to interpret and respond to human mental states is crucial for enabling effective interactions. In this paper, we review LLMs’ ToM capabilities by analyzing both evaluation benchmarks and enhancement strategies. For evaluation, we focus on recently proposed and widely used story-based benchmarks. For enhancement, we provide an in-depth analysis of recent methods aimed at improving LLMs’ ToM abilities. Furthermore, we outline promising directions for future research to further advance these capabilities and better adapt LLMs to more realistic and diverse scenarios. Our survey serves as a valuable resource for researchers interested in evaluating and advancing LLMs’ ToM capabilities.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have achieved remarkable success across various tasks, particularly in Natural Language Processing (NLP) Qin et al. (2023). Currently, there is growing interest to explore the cognitive abilities of LLMs through novel and intriguing tasks, including assessments of their Emotional Intelligence (EI) Yang et al. (2023); Huang et al. (2024); Amin et al. (2024); Sabour et al. (2024); Liu et al. (2024), strategic reasoning capabilities Gandhi et al. (2023b); Zhang et al. (2024b), and ToM capabilities Wu et al. (2023); van Duijn et al. (2023); Wu et al. (2024); Shapira et al. (2024); Chen et al. (2024); Jin et al. (2024). Among these capabilities, there is an ongoing debate about whether LLMs truly possess ToM abilities. Some studies suggest that LLMs demonstrate promising signs of ToM competence Bubeck et al. (2023); Kosinski (2024); Street et al. (2024), while others contend that these abilities are often superficial and unstable Shapira et al. (2024); Ullman (2023); Ma et al. (2023a); Verma et al. (2024); Zhu et al. (2024); Bortoletto et al. (2024a).
Figure 1: This paper reviews both the evaluation and enhancement of theory of mind capabilities in LLMs. For evaluation, we cover passive, story-based benchmarks and adaptable interactive benchmarks. For enhancement, we discuss recent effective methods, including prompt-only approaches and strategies incorporating additional techniques, as well as earlier methods from the era of small language models. Due to space constraints, interactive benchmarks and earlier methods (indicated in the dashed area) are provided in the Appendix E and Appendix F.
@@
Consequently, as illustrated in Figure 1, our survey covers both the evaluation and enhancement of LLMs’ ToM capabilities. Specifically, we highlight recently proposed and widely used story-based benchmarks from 2023–2024 that have not been covered in existing reviews, analyzing their evolution over time. In addition, we provide a detailed review of the latest methods designed to improve ToM performance in LLMs, aiming to lay a foundation for future research in this area. To the best of our knowledge, this is the first study to offer such an in-depth analysis of both the assessment and advancement of ToM capabilities in LLMs. Our main contributions are summarized as follows:
  • Broad Survey. This paper provides the first broad survey that addresses both the evaluation and enhancement of LLMs’ ToM capabilities. It offers a detailed analysis of recently proposed story-based benchmarks and the strategies developed to improve LLM performance in this domain.
@@ 2 Theory of Mind
The term "orders" refers to the number of mental state attributions required to answer a given question Wu et al. (2023). For example, the question "Where will Sally look for her marble?" is considered first-order, while "Where does Anne think Sally will look for her marble?" is classified as second-order. When Sally’s belief does not align with the reality, the question is categorized as a false-belief question; conversely, if her belief is accurate, it is referred to as a true-belief question.
@@
As shown in Table 1, more than ten benchmarks have been recently proposed to assess the ToM capabilities of LLMs. The majority of these benchmarks are based on well-established psychological tests, including the Sally-Anne Test Baron-Cohen et al. (1985) and the Smarties Task Perner et al. (1987); Gopnik and Astington (1988), both of which are discussed in detail in Appendix A.
@@
Benchmarks Mental States
ToMi Le et al. (2019)
MindGames Sileo and Lernould (2023)
HI-TOM Wu et al. (2023)
TOMCHALLENGES Ma et al. (2023a)
FANTOM Kim et al. (2023)
BigToM Gandhi et al. (2023a)
OpenToM Xu et al. (2024a)
NegotiationToM Chan et al. (2024)
TOMBENCH Chen et al. (2024)
COMMON-TOM Soubki et al. (2024)
SimpleToM Gu et al. (2024)
MMToM-QA Jin et al. (2024)
MuMA-ToM Shi et al. (2025)
Table 1: Benchmarks and their coverage of mental states. The mental state categories are derived from the ATOMS Beaudoin et al. (2020). In this paper, we treat the "goals" referenced in MMToM-QA Jin et al. (2024) and MuMA-ToM Shi et al. (2025) as equivalent to "intentions" in ATOMS. Benchmarks are divided into two categories: text-only benchmark and multimodal benchmark.
@@ 3 Benchmarks for Evaluating ToM Capabilities in LLMs
3.1 Text-Only Benchmarks
@@ 3.1.1 Overview of Text-Only Benchmarks
TOMCHALLENGES Ma et al. (2023a). Ma et al. (2023a) identified that inconsistencies in evaluating LLMs’ ToM capabilities might arise from variations in evaluation methods and prompts. To address this, they designed six distinct questions for each "order" of ToM reasoning about each narrative story, using templates that span three generation settings: fully-constrained, semi-constrained, and open-ended. The stories are adapted from classic tasks like the Sally-Anne test and Smarties tasks Baron-Cohen et al. (1985); Perner et al. (1987); Gopnik and Astington (1988), covering up to second-order belief questions. To enhance grading efficiency, they proposed an auto-grader based on the GPT-4222 https://platform.openai.com/docs/models/gpt-4 model. However, the effectiveness of this automatic evaluation tool may depend on the model’s own ToM capabilities.
@@
NegotiationToM Chan et al. (2024). Based on the Belief-Desire-Intention (BDI) framework Bratman (1987), NegotiationToM evaluates machine ToM in multi-round real-world negotiation dialogues, assessing mental states like desires, second-order beliefs, and intentions. Built on the CaSiNo dataset Chawla et al. (2021), it is the first ToM benchmark in negotiation contexts but remains passive and may suffer from data contamination due to reuse of existing data.
@@ 3.1.2 Trends in Text-Only Benchmarks
  • Questions. To enable robust evaluation, benchmark questions have become more diverse and complex, evolving from single question answering taskLe et al. (2019) to multiple question answering tasksMa et al. (2023a); Kim et al. (2023). Additionally, the content of the questions has changed, for instance, asking about the beliefs of a group of people rather than a single individual Sileo and Lernould (2023); Kim et al. (2023).
  • Mental States. Benchmarks have expanded from evaluating only beliefs Le et al. (2019) to covering a wider range of mental states Chan et al. (2024); Chen et al. (2024), including percepts, desires, and even social commonsense knowledge Xu et al. (2024a).
Approaches Order Tested Fine-Tuning
SYMBOLICTOM Sclar et al. (2023) Up to third-order No
SIMTOM Wilf et al. (2024) Up to second-order No
PercepToM Jung et al. (2024) Up to second-order No
TIMETOM Hou et al. (2024) Up to third-order No
ToM-LM Tang and Belle (2024) Up to second-order Yes
BIP-ALM Jin et al. (2024) Only first-order Yes
LIMP Shi et al. (2025) Up to second-order No
3.2 Multimodal Benchmarks
In addition to the aforementioned developments, some researchers have recognized that ToM reasoning goes beyond simple text comprehension, which has led to the introduction of several multimodal ToM benchmarks.
@@ 3.2 Multimodal Benchmarks
MuMA-ToM Shi et al. (2025). Building on videos generated in VirtualHome Puig et al. (2018), this benchmark introduces the first multimodal, multi-agent ToM evaluation focusing on mental reasoning in embodied interactions. It includes three types of questions: belief inference, social goal inference, and belief-of-goal inference (a form of second-order belief question). Each question presents three options, with one being the most likely to be correct. Unlike MMToM-QA Jin et al. (2024), MuMA-ToM provides distinctly separate information available through specific modalities, enabling a deeper understanding of how models integrate multimodal inputs to accurately answer each question. The text input is presented in two forms: conversations between two agents in the video or textual descriptions of half of the video. MuMA-ToM is currently limited to three social goals and two agents. Additionally, as it relies on video content from VirtualHome, further efforts are required to enhance its sim-to-real transferability for real-world applications.
@@
Figure 2: A comparison of methods for enhancing LLMs’ ToM capabilities through different prompting techniques, using the following narrative as context. Story: Alice and Bob are in a room with a basket and a box. Alice puts some celery in the basket and leaves the room. Bob then moves the celery into the box Sclar et al. (2023).
@@ 5 Future Directions
Most evaluations indicate that LLMs still lack robust ToM reasoning abilities. To address this, future work should focus on developing better strategies and building more datasets—both for evaluation and fine-tuning. In this section, we first outline shared future directions for benchmarks and enhancement strategies, then discuss those specific to either goal [^4] Due to space limitations, additional future directions are discussed in the Appendix G.. Common future directions for both benchmarks and strategies involve:
  • Expanding the scope of mental states. As shown in Table 1 and Table 2, current benchmarks and strategies mainly focus on belief-related reasoning, while other mental states require more attention Ma et al. (2023b). Future benchmarks should explore a broader range of mental states. To better evaluate LLMs’ ToM capabilities, a wider variety of psychological tests should be integrated into the benchmarking process. Fu et al. (2023) compiled a list of 127 tests designed to measure the ToM capabilities of children from birth to 12 years old, which could be further adapted for testing LLMs as the technology advances.
  • Addressing multi-modal ToM reasoning. Humans interact with the world through multiple channels—an aspect that simple text-based stories cannot fully capture Liu et al. (2023). Research on ToM capabilities in LLMs aims to develop models that function as agents with robust ToM skills, potentially achieving human-level competence Ma et al. (2023b). Achieving this goal requires the ability to infer mental states from visual, auditory, contextual, and other cues. For instance, incorporating multimodal content such as short films Dziobek et al. (2006) or cartoons Völlm et al. (2006); Parmar et al. (2024) into evaluation benchmarks could enhance ToM assessments. Moreover, current methods primarily focus on multiple-choice question answering in household contexts, and converting video information to text may lead to the loss of crucial details. Developing strategies that effectively handle multimodal inputs and preserve comprehensive information remains an important direction for future research.
  • Active benchmarks and strategies for agentic decision-making. Passive benchmarks Ma et al. (2023b); Chan et al. (2024) are insufficient. More research is needed to leverage LLMs as agents capable of making decisions in complex environments Ma et al. (2023b); Li et al. (2023); Zhou et al. (2023a, b), enabling deeper investigation of their ToM capabilities.
@@
  • Exploring joint approaches. Most existing methods rely on pipeline architectures with no feedback loops between stages. Error propagation is a significant issue in pipeline approaches Yang and Mitchell (2016); Liu et al. (2018). For example, in SYMBOLICTOM, if the belief graph for a character is incorrect, it becomes difficult to obtain the correct answer when the question relates to this flawed belief graph. However, if conflicts are detected during the inference stage and feedback is provided during the belief graph construction stage, the accuracy of the inference and the coherence of the results can be improved. Investigating joint or iterative approaches that integrate feedback to improve ToM reasoning remains an open research question.
  • Progressive learning strategies for tackling ToM task complexity. Given that ToM tasks can vary significantly in complexity, a continual/curriculum learning Chen and Liu (2018); Soviany et al. (2022); Wang et al. (2024) strategy may be needed to address these challenges progressively, starting with simpler tasks and advancing to more complex ones.
Finally, we believe that evaluating reasoning processes is essential. Assessing answer correctness alone is insufficient—datasets should be designed to support reasoning evaluation Kawabata and Sugawara (2023); Jung et al. (2024); Xu et al. (2024b), calling for more effective and automated evaluation strategies. Likewise, probing a model’s internal representations of mental states Bortoletto et al. (2024b) is an intriguing direction that warrants further exploration.
@@ 6 Conclusion
Acknowledgments
We sincerely thank all reviewers for their valuable feedback. This work was supported by A*STAR CRF funding awarded to Cheston Tan.
Appendix A Psychological Tests
@@ A.1 Sally-Anne Test
  • Reality Question: "Where is the marble really?"
  • Memory Question: "Where was the marble in the beginning?"
@@ A.3 Smarties Task
The Smarties task Perner et al. (1987), sometimes referred to as "unexpected contents" [^5] https://en.wikipedia.org/wiki/Theory_of_mind or "appearance-reality" [^6] http://www.sfu.ca/~hedberg/ToM.pdf task, involves asking children what they think is inside a box that looks like it contains smarties. After the child guesses "smarties", they are shown that the box actually holds a different item, such as pencils. The box is then resealed, and the child is asked what they think someone else, who hasn’t seen the real contents, will believe is inside. The child succeeds if they say the other person will think "smarties" are in the box, and fails if they say the other person will think it contains pencils. According to Gopnik and Astington’s research, children generally pass this task by the age of four or five Gopnik and Astington (1988).